The problem is
1.2x-3.5=x+1.7
the other is
6-7(m-3)+1=0

Answer by checkley77(12844)

1.2x-3.5=x+1.7
1.2x-x=1.7+3.5
.2x=5.2
x=5.2/.2
x=26 answer.
------------------------------
6-7(m-3)+1=0
6-7m+21+1=0
-7m=-6-21-1
-7m=-28
m=-28/-7
m=4 answer.